OpenStudy (anonymous):

help:)) separate 114 into two parts such that when the larger is divided by the smaller , the quotient is 3 and the remainder is 18..?

OpenStudy (anonymous):

i mean 90 and 24

OpenStudy (anonymous):

HAHAHA. 90/24

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

help:(

OpenStudy (anonymous):

I gave you the answer. mr.jimmyrep gave you the solution

OpenStudy (anonymous):

sorry i gotta go

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

sorry i totally dont get the answer

OpenStudy (anonymous):

let the 2 numbers be 114-x and x (114 - x )/ x gives a quotient of 3 and a remainder of 18 so the equation becomes: 114-x / x = 3 + 18/ x multiply by x 114 - x = 3x + 18 4x = 96 x = 24 so the 2 numbers are 90 and 24